# Solving Optimal Experiment Design with Mixed-Integer Convex Methods

### v6

@@ 1 Introduction
The problem is running all experiments, potentially even multiple times to account for errors, is often not realistic because of time and cost constraints. Thus, OEDP deals with finding a subset of size $N$ of the experiments providing the ‘‘most information’’ about the experiment space (Pukelsheim, 2006; de Aguiar et al., 1995). In general, the number of allowed experiments $N$ is in the interval $[n,m]$ in order to allow a solution to the regression model in Equation 1.
In Section 2, we investigate the necessary conditions for a function to be a valid and useful information measure. Every information function leads to a different criterion. In this paper, we focus on two popular criteria, namely the A-criterion and D-criterion, see Section 2.1.
In general, OEDP leads to a Mixed-Integer Non-Linear Problem (MINLP). There has been a lot of development in the last years in solving MINLP (Kronqvist et al., 2019). Nevertheless, the capabilities of current MINLP solvers are far away from their linear counterparts, the Mixed-Integer Problem (MIP) solvers (Bestuzheva et al., 2021), especially concerning the magnitude of the problems that can be handled. Therefore, instead of solving the actual MINLP formulation, a continuous version of OEDP, called the Limit Problem, is often solved and the integer solution is created from the continuous solution by rounding schemes and other heuristics (Pukelsheim, 2006). This does not necessarily lead to optimal solutions, though, and the procedure is not always applicable to a given continuous solution either because it requires a certain sparsity of the continuous solution.
The goal of this paper is to compare the performance of different MINLP approaches for OEDP problems. A special focus is put on the newly proposed framework Boscia (Hendrych et al., 2025b) which can solve larger instances and significantly outperforms the other examined approaches. The new method leverages Frank-Wolfe algorithms and a formulation different from the other approaches. In Section 3, we establish the convergence of the Frank-Wolfe algorithm on the continuous relaxations of the considered OEDP problems. The different solution methods are detailed in Section 4 and evaluation of the computational experiments can be found in Section 5.
@@ 1.1 Related work
As mentioned, one established method of solution is the reduction to a simpler problem by removing the integer constraints and employing heuristics to generate an integer solution from the continuous solution. Recently, there have been more publications tackling the MINLP formulation of the Optimal Experiment Design Problem. These, however, concentrate on specific information measures, in particular the A-criterion, see (Nikolov et al., 2022; Ahipaşaoğlu, 2015), and the D-criterion, (Welch, 1982; Ponte et al., 2025; Li et al., 2024; Xu & Liberti, 2024; Ponte et al., 2023; Fampa & Lee, 2022; Chen et al., 2024; Ponte et al., 2024). The most general solution approach known to the authors was introduced in Ahipaşaoğlu (2021). It considers OEDP under matrix means which, in particular, includes the A-criterion and D-criterion. While the matrix means covers many information measures of interest, it still yields a restricting class of information functions. For example, the G-criterion and V-criterion are not included in this class of functions (de Aguiar et al., 1995). The newly proposed framework Boscia only requires the information measures to be either $L$-smooth, i.e. the gradient is Lipschitz continuous, or generalized self-concordant, thereby covering a larger group of information functions. In addition, Boscia does not suppose any prior knowledge about the structure of the problem, being thus more flexible in terms of problem formulations. On the other hand, it is highly customizable, giving the user the ability to exploit the properties of their problems to speed up the solving process. An in-depth and unified theory for the Optimal Experiment Design Problem can be found in Pukelsheim (2006).
@@ 1.2 Contribution
First, we propose a unified view of multiple experiment design formulations as the optimization of a nonlinear (not necessarily Lipschitz-smooth) information function over a truncated scaled probability simplex intersected with the set of integers. Unlike most other formulations that replace the nonlinear objective with nonlinear and/or conic constraints, we preserve the original structure of the problem. Additionally, we can easily handle special cases of OEDP without any reformulations and additional constraints on OEDP since we do not suppose a specific problem structure, unlike the approach in Ahipaşaoğlu (2021).
@@
Finally, we use the recently proposed Boscia (Hendrych et al., 2025b) framework to solve the integer optimization problem with a Frank-Wolfe method for the node relaxations of the Branch-and-Bound tree and show the effectiveness of the method on instances generated with various degrees of correlation between the parameters.
@@ 1.3 Notation
In the following let $\lambda_{i}(A)$ denote the $i$-th eigenvalue of matrix $A$; we assume that these are sorted in increasing order. Moreover, $\lambda_{\min}(A)$ and $\lambda_{\max}(A)$ denote the minimum and maximum eigenvalue of $A$, respectively. We define and denote the condition number of a positive semi-definite matrix $A$ as $\kappa(A)=\frac{\lambda_{\max}(A)}{\lambda_{\min}(A)}$. Further, let $\log\det(\cdot)$ be the log-determinant of a positive definite matrix. Given matrices $A$ and $B$ of same dimensions, $A\circ B$ denotes their Hadamard product. Given a vector $\mathbf{x}$, $\operatorname{diag}(\mathbf{x})$ denotes the diagonal matrix with $\mathbf{x}$ on its diagonal and let $\operatorname{Diag}(A)$ denote the diagonal of the matrix $A$. The cones of positive definite and positive semi-definite matrices in ${\mathbb{R}}^{n\times n}$ will be denoted by $\mathbb{S}_{++}^{n}$ and $\mathbb{S}_{+}^{n}$, respectively. We will refer to them as PD and PSD cones. The set of symmetric $n\times n$ matrices will be denoted by $\mathbb{S}^{n}$. Let ${\mathbb{N}}_{0}$ denote the set of positive integers including 0 and for $m\in{\mathbb{N}}$ let $[m]=\{1,2,\dots,m\}$. Lastly, we denote matrices with capital letters, e.g. $X$, vectors with bold small letters, e.g. $\mathbf{x}$, and simple small letters are scalars, e.g. $\lambda$.
@@ 2 The optimal experiment design problem
It is a measure of the variance of the experiment parameters (Ahipaşaoğlu, 2015). Our goal is to obtain the ‘‘most’’ information, i.e. we want to maximize some measure of the information matrix. This is equivalent to minimizing over the dispersion matrix (Pukelsheim, 2006), thereby minimizing the variance of the parameters.
@@
Definition 2.1 (Information Function (Pukelsheim, 2006)).
@@
The most frequently used information functions arise from matrix means $\phi_{p}$ (Pukelsheim (2006), Ahipaşaoğlu (2015)). The matrix means are defined as follows:
Definition 2.2 (Matrix mean).
@@
Note that the function $\phi_{p}$ satisfies the requirements of Definition 2.1 only for $p\leq 1$ (Pukelsheim, 2006). The different values of $p$ lead to different so-called criteria. The Optimal Experiment Design Problem using a matrix means function is then defined as
@@
Using the logarithm formulation, one can recover Fenchel duality results (Sagnol, 2010, Chapter 3). It can also by beneficial from a information theory point of view. Note, however, that the $\log$ does not necessarily preserve concavity.
The resulting problem (OEDP) is an integer optimization problem which, depending on the information function $\phi_{p}$, can be $\mathcal{NP}$-hard. The two information measures we will focus on lead to $\mathcal{NP}$-hard problems (Welch, 1982; Nikolov et al., 2022; Li, 2025; Li et al., 2024). Thus, these OEDP are hard to solve. Often, therefore, one solves the so-called Limit Problem obtained by letting the number of allowed experiments $N$ go to infinity. The resulting optimization problem is continuous.
@@
The variable vector $\mathbf{w}$ can be interpreted as a probability distribution. Under suitable assumptions, one can generate a finite experiment design from the solution of the limit problem, see Ahipaşaoğlu (2015), Pukelsheim (2006, Chapter 12). One assumption is that the support of the solution of the Limit Problem (OEDP-Limit) is smaller than the allowed number of experiments $N$. This is often not the case (Pukelsheim, 2006). If an integer solution can be obtained, it is feasible but not necessarily optimal. Optimality can be achieved in some special cases, see Pukelsheim (2006). However, these have very specific requirements that are often not achievable.
@@ 2.1 The A-, D- and GTI-optimal experiment design problem
For this paper, we focus on two particular criteria arising from the matrix means. The two most commonly used criteria are the D-optimality and A-optimality criterion, $p=0$ and $p=-1$, respectively.
@@
Choosing $p=0$ in Definition 2.2, yields for the objective of (OEDP)
@@
Observe that it is equivalent to minimize $\log\det(X(\mathbf{x})^{-1})=\log\det D(\mathbf{x})$, so the determinant of the dispersion matrix. This is also called the generalized variance of the parameter $\bm{\theta}$ (Pukelsheim, 2006). Thus, a maximal value of $\det X$ corresponds to a maximal volume of standard ellipsoidal confidence region of $\bm{\theta}$ (Ponte et al., 2025). Additionally, the D-criterion is invariant under reparameterization (Pukelsheim, 2006).
@@
For parameters with a physical interpretation, the A-optimality criterion is a good choice (Pukelsheim, 2006) as it amounts to minimizing the average of the variances of $\bm{\theta}$. Geometrically, the A-criterion amounts to minimizing the diagonal of the bounding box of the ellipsoidal confidence region (Sagnol, 2010). The resulting optimization problem can be stated in two different ways, keeping the $\log$ and using the $\log$ rules and the positive homogeneity of the information functions.
@@ 3 Properties of the problems and convergence guarantees
To ensure that the newly proposed framework in Hendrych et al. (2025b) is applicable to the previously introduced problems, we have to show that the Frank-Wolfe algorithms used as the subproblem solvers converge on the continuous subproblems. Frank-Wolfe normally requires the objective function to be convex and $L$-smooth. The convexity part is trivial, for proofs see Section A.2. The $L$-smoothness property only holds for the Fusion Problems on the whole feasible region. For the Optimal Problems, we will show local $L$-smoothness in Section 3.1 and an alternative property guaranteeing convergence of Frank-Wolfe, namely (generalized) self-concordance, in Section 3.2.
Lastly, we establish sharpness or the Hölder error bound condition of the objective functions, therefore yielding improved convergence rates for Frank-Wolfe methods compared to the standard $\mathcal{O}(1/\epsilon)$ rate where $\epsilon$ is the additive error in the primal gap, see Section 3.3.
@@ 3.1 Lipschitz smoothness
In Section A.3, the reader can find $L$-smoothness proofs for $f(X)$, $g(X)$ and $k(X)$ over
@@
Observe that the Hessian expressions are very similar. Hence, we will only present the proof for $g_{F}$. The proof of $f_{F}$ follows the same argumentation. By Horn (1990, page 95) it is known that the maximum eigenvalue of a Hadamard product between two positive semi-definite matrices $N$ and $M$, $\lambda_{\max}(N\circ M)$, is upper bounded by the maximum diagonal entry of $N$ and the largest eigenvalue of $M$, so
@@
In our case, $N=AA^{\intercal}$ and $M=AX_{C}^{-p-2}A^{\intercal}$. By Corollary A.19 in Section A.1, we have that $X_{C}^{-p-2}$ is positive definite. Thus, both factors are semi-definite and we may use the provided bound.
@@
By Lemma A.18, we have $\lambda(A^{p})=\lambda(A)^{p}$.
@@
By the definition of $X$ and by Weyl’s inequality on the eigenvalues of sums of positive semi-definite matrices Weyl (1912), we have that
@@
As before, we omit the dependency on $\mathbf{x}$ for better readability. Note that both parts of the nominator are symmetric matrices. Thus, we can use the upper bound on the eigenvalues of sums of symmetric matrices in Lemma A.28. Further, note that for any $\mathbf{z}\in{\mathbb{R}}^{n}$ the matrix $\mathbf{z}\mathbf{z}^{\intercal}$ is PSD and its maximum eigenvalue is $\left\langle\mathbf{z},\mathbf{z}\right\rangle$.
@@
An upper bound on the maximum eigenvalue of $\left(AA^{\intercal}\right)\circ\left(AX_{C}^{-p-2}A^{\intercal}\right)$ is already provided by Theorem 3.4. Observe that
@@
As stated before, the proofs of Theorems 3.4 and 3.5 work because we can bound the minimum eigenvalue of the Fusion information matrix $X_{C}$ away from zero thanks to the regular matrix $C$. This is not the case of the information matrices $X(\mathbf{x})$ of the Optimal Problems as they can be singular for some $x\in\mathcal{P}$. Therefore, the objectives of the Optimal Problems are not $L$-smooth over $\mathcal{P}$. We can show, however, that the objectives are $L$-smooth on a restricted area of the feasible region.
@@
where ()(*)( ∗ ) is a placeholder for $f(\mathbf{x})$, $g(\mathbf{x})$ and $k(\mathbf{x})$, respectively. Using the bound provided by Merikoski & Virtanen (1997, Theorem 1) and the fact that the objective value is decreasing, we can bound the minimum eigenvalue by quantities depending on the experiment matrix $A$ and the start point $\mathbf{x}_{0}$. Let further $X_{0}=X(\mathbf{x}_{0})$.
Theorem 3.6.
@@
To bound the Lipschitz constant, we need lower bound the minimum eigenvalue of $X(\mathbf{x})$. We start with $L_{f}$ and find by using the bound provided by Merikoski & Virtanen (1997, Theorem 1)
@@
The trace can be upper bounded by $\operatorname{Tr}(X)\leq mN\max_{i}\left\|\mathbf{v}_{i}\right\|_{2}^{2}$ and by knowing that $-\log\det(X(\mathbf{x}))\leq-\log\det(X_{0})$, we find

### v7

@@ 1 Introduction
The problem is running all experiments, potentially even multiple times to account for errors, is often not realistic because of time and cost constraints. Thus, OEDP deals with finding a subset of size $N$ of the experiments providing the “most information” about the experiment space (pukelsheim2006optimal; de1995d). In general, the number of allowed experiments $N$ is in the interval $[n,m]$ in order to allow a solution to the regression model in Equation˜1.
In Section˜2, we investigate the necessary conditions for a function to be a valid and useful information measure. Every information function leads to a different criterion. In this paper, we focus on two popular criteria, namely the A-criterion and D-criterion, see Section˜2.1.
In general, OEDP leads to a Mixed-Integer Non-Linear Problem (MINLP). There has been a lot of development in the last years in solving MINLP (kronqvist2019review). Nevertheless, the capabilities of current MINLP solvers are far away from their linear counterparts, the Mixed-Integer Problem (MIP) solvers (bestuzheva2021scip), especially concerning the magnitude of the problems that can be handled. Therefore, instead of solving the actual MINLP formulation, a continuous version of OEDP, called the Limit Problem, is often solved and the integer solution is created from the continuous solution by rounding schemes and other heuristics (pukelsheim2006optimal). This does not necessarily lead to optimal solutions, though, and the procedure is not always applicable to a given continuous solution either because it requires a certain sparsity of the continuous solution.
The goal of this paper is to compare the performance of different MINLP approaches for OEDP problems. A special focus is put on the newly proposed framework Boscia (hendrych2023convex) which can solve larger instances and significantly outperforms the other examined approaches. The new method leverages Frank-Wolfe algorithms and a formulation different from the other approaches. In Section˜3, we establish the convergence of the Frank-Wolfe algorithm on the continuous relaxations of the considered OEDP problems. The different solution methods are detailed in Section˜4 and evaluation of the computational experiments can be found in Section˜5.
@@ 1.1 Related work
As mentioned, one established method of solution is the reduction to a simpler problem by removing the integer constraints and employing heuristics to generate an integer solution from the continuous solution. Recently, there have been more publications tackling the MINLP formulation of the Optimal Experiment Design Problem. These, however, concentrate on specific information measures, in particular the A-criterion, see (nikolov2022proportional; ahipacsaouglu2015first), and the D-criterion, (welch1982branch; ponte2023branch; li2022d; xu2024submodular; ponte2023computing; fampa2022maximum; chen2024generalized; ponte2024convex). The most general solution approach known to the authors was introduced in ahipacsaouglu2021branch. It considers OEDP under matrix means which, in particular, includes the A-criterion and D-criterion. While the matrix means covers many information measures of interest, it still yields a restricting class of information functions. For example, the G-criterion and V-criterion are not included in this class of functions (de1995d). The newly proposed framework Boscia only requires the information measures to be either $L$-smooth, i.e. the gradient is Lipschitz continuous, or generalized self-concordant, thereby covering a larger group of information functions. In addition, Boscia does not suppose any prior knowledge about the structure of the problem, being thus more flexible in terms of problem formulations. On the other hand, it is highly customizable, giving the user the ability to exploit the properties of their problems to speed up the solving process. An in-depth and unified theory for the Optimal Experiment Design Problem can be found in pukelsheim2006optimal.
@@ 1.2 Contribution
First, we propose a unified view of multiple experiment design formulations as the optimization of a nonlinear (not necessarily Lipschitz-smooth) information function over a truncated scaled probability simplex intersected with the set of integers. Unlike most other formulations that replace the nonlinear objective with nonlinear and/or conic constraints, we preserve the original structure of the problem. Additionally, we can easily handle special cases of OEDP without any reformulations and additional constraints on OEDP since we do not suppose a specific problem structure, unlike the approach in ahipacsaouglu2021branch.
@@
Finally, we use the recently proposed Boscia (hendrych2023convex) framework to solve the integer optimization problem with a Frank-Wolfe method for the node relaxations of the Branch-and-Bound tree and show the effectiveness of the method on instances generated with various degrees of correlation between the parameters.
@@ 1.3 Notation
In the following let $\lambda_{i}(A)$ denote the $i$-th eigenvalue of matrix $A$; we assume that these are sorted in increasing order. Moreover, $\lambda_{\min}(A)$ and $\lambda_{\max}(A)$ denote the minimum and maximum eigenvalue of $A$, respectively. We define and denote the condition number of a positive semi-definite matrix $A$ as $\kappa(A)=\frac{\lambda_{\max}(A)}{\lambda_{\min}(A)}$. Further, let $\log\det(\cdot)$ be the log-determinant of a positive definite matrix. Given matrices $A$ and $B$ of same dimensions, $A\circ B$ denotes their Hadamard product. Given a vector $\mathbf{x}$, $\operatorname{diag}(\mathbf{x})$ denotes the diagonal matrix with $\mathbf{x}$ on its diagonal and let $\operatorname{Diag}(A)$ denote the diagonal of the matrix $A$. The cones of positive definite and positive semi-definite matrices in ${\mathbb{R}}^{n\times n}$ will be denoted by $\mathbb{S}_{++}^{n}$ and $\mathbb{S}_{+}^{n}$, respectively. We will refer to them as PD and PSD cones. The set of symmetric $n\times n$ matrices will be denoted by $\mathbb{S}^{n}$. Let ${\mathbb{N}}_{0}$ denote the set of positive integers including 0 and for $m\in{\mathbb{N}}$ let $m]=\{1,2,\dots,m\}$. Lastly, we denote matrices with capital letters, e.g. $X$, vectors with bold small letters, e.g. $\mathbf{x}$, and simple small letters are scalars, e.g. $\lambda$.
@@ 2 The optimal experiment design problem
It is a measure of the variance of the experiment parameters (ahipacsaouglu2015first). Our goal is to obtain the “most” information, i.e. we want to maximize some measure of the information matrix. This is equivalent to minimizing over the dispersion matrix (pukelsheim2006optimal), thereby minimizing the variance of the parameters.
@@
Definition 1 (Information Function (pukelsheim2006optimal)).
@@
The most frequently used information functions arise from matrix means $\phi_{p}$ (pukelsheim2006optimal, ahipacsaouglu2015first). The matrix means are defined as follows:
Definition 2 (Matrix mean).
@@
Note that the function $\phi_{p}$ satisfies the requirements of Definition˜1 only for $p\leq 1$ (pukelsheim2006optimal). The different values of $p$ lead to different so-called criteria. The Optimal Experiment Design Problem using a matrix means function is then defined as
@@
Using the logarithm formulation, one can recover Fenchel duality results (sagnol2010optimal, Chapter 3). It can also by beneficial from a information theory point of view. Note, however, that the $\log$ does not necessarily preserve concavity.
The resulting problem (OEDP) is an integer optimization problem which, depending on the information function $\phi_{p}$, can be $\mathcal{NP}$-hard. The two information measures we will focus on lead to $\mathcal{NP}$-hard problems (welch1982branch; nikolov2022proportional; li2025strong; li2022d). Thus, these OEDP are hard to solve. Often, therefore, one solves the so-called Limit Problem obtained by letting the number of allowed experiments $N$ go to infinity. The resulting optimization problem is continuous.
@@
The variable vector $\mathbf{w}$ can be interpreted as a probability distribution. Under suitable assumptions, one can generate a finite experiment design from the solution of the limit problem, see ahipacsaouglu2015first, pukelsheim2006optimal. One assumption is that the support of the solution of the Limit Problem (OEDP-Limit) is smaller than the allowed number of experiments $N$. This is often not the case (pukelsheim2006optimal). If an integer solution can be obtained, it is feasible but not necessarily optimal. Optimality can be achieved in some special cases, see pukelsheim2006optimal. However, these have very specific requirements that are often not achievable.
@@ 2.1 The A-, D- and GTI-optimal experiment design problem
For this paper, we focus on two particular criteria arising from the matrix means. The two most commonly used criteria are the D-optimality and A-optimality criterion, $p=0$ and p=1p=-1, respectively.
@@
Choosing $p=0$ in Definition˜2, yields for the objective of (OEDP)
@@
Observe that it is equivalent to minimize $\log\det(X(\mathbf{x})^{-1})=\log\det D(\mathbf{x})$, so the determinant of the dispersion matrix. This is also called the generalized variance of the parameter $\bm{\theta}$ (pukelsheim2006optimal). Thus, a maximal value of $\det X$ corresponds to a maximal volume of standard ellipsoidal confidence region of $\bm{\theta}$ (ponte2023branch). Additionally, the D-criterion is invariant under reparameterization (pukelsheim2006optimal).
@@
For parameters with a physical interpretation, the A-optimality criterion is a good choice (pukelsheim2006optimal) as it amounts to minimizing the average of the variances of $\bm{\theta}$. Geometrically, the A-criterion amounts to minimizing the diagonal of the bounding box of the ellipsoidal confidence region (sagnol2010optimal). The resulting optimization problem can be stated in two different ways, keeping the $\log$ and using the $\log$ rules and the positive homogeneity of the information functions.
@@ 3 Properties of the problems and convergence guarantees
To ensure that the newly proposed framework in hendrych2023convex is applicable to the previously introduced problems, we have to show that the Frank-Wolfe algorithms used as the subproblem solvers converge on the continuous subproblems. Frank-Wolfe normally requires the objective function to be convex and $L$-smooth. The convexity part is trivial, for proofs see Section˜A.2. The $L$-smoothness property only holds for the Fusion Problems on the whole feasible region. For the Optimal Problems, we will show local $L$-smoothness in Section˜3.1 and an alternative property guaranteeing convergence of Frank-Wolfe, namely (generalized) self-concordance, in Section˜3.2.
Lastly, we establish sharpness or the Hölder error bound condition of the objective functions, therefore yielding improved convergence rates for Frank-Wolfe methods compared to the standard $\mathcal{O}(1/\epsilon)$ rate where $\epsilon$ is the additive error in the primal gap, see Section˜3.3.
@@ 3.1 Lipschitz smoothness
In Section˜A.3, the reader can find $L$-smoothness proofs for $f(X)$, $g(X)$ and $k(X)$ over
@@
Observe that the Hessian expressions are very similar. Hence, we will only present the proof for $g_{F}$. The proof of $f_{F}$ follows the same argumentation. By horn1990hadamard it is known that the maximum eigenvalue of a Hadamard product between two positive semi-definite matrices $N$ and $M$, $\lambda_{\max}(N\circ M)$, is upper bounded by the maximum diagonal entry of $N$ and the largest eigenvalue of $M$, so
@@
In our case, $N=AA^{\intercal}$ and $M=AX_{C}^{-p-2}A^{\intercal}$. By Corollary˜2 in Section˜A.1, we have that $X_{C}^{-p-2}$ is positive definite. Thus, both factors are semi-definite and we may use the provided bound.
@@
By Lemma˜6, we have $\lambda(A^{p})=\lambda(A)^{p}$.
@@
By the definition of $X$ and by Weyl’s inequality on the eigenvalues of sums of positive semi-definite matrices weyl1912asymptotische, we have that
@@
As before, we omit the dependency on $\mathbf{x}$ for better readability. Note that both parts of the nominator are symmetric matrices. Thus, we can use the upper bound on the eigenvalues of sums of symmetric matrices in Lemma˜15. Further, note that for any $\mathbf{z}\in{\mathbb{R}}^{n}$ the matrix $\mathbf{z}\mathbf{z}^{\intercal}$ is PSD and its maximum eigenvalue is $\left\langle\mathbf{z},\mathbf{z}\right\rangle$.
@@
An upper bound on the maximum eigenvalue of $\left(AA^{\intercal}\right)\circ\left(AX_{C}^{-p-2}A^{\intercal}\right)$ is already provided by Theorem˜1. Observe that
@@
As stated before, the proofs of Theorems˜1 and 2 work because we can bound the minimum eigenvalue of the Fusion information matrix $X_{C}$ away from zero thanks to the regular matrix $C$. This is not the case of the information matrices $X(\mathbf{x})$ of the Optimal Problems as they can be singular for some $x\in\mathcal{P}$. Therefore, the objectives of the Optimal Problems are not $L$-smooth over $\mathcal{P}$. We can show, however, that the objectives are $L$-smooth on a restricted area of the feasible region.
@@
where ()(*) is a placeholder for $f(\mathbf{x})$, $g(\mathbf{x})$ and $k(\mathbf{x})$, respectively. Using the bound provided by merikoski1997bounds and the fact that the objective value is decreasing, we can bound the minimum eigenvalue by quantities depending on the experiment matrix $A$ and the start point $\mathbf{x}_{0}$. Let further $X_{0}=X(\mathbf{x}_{0})$.
Theorem 3.
@@
To bound the Lipschitz constant, we need lower bound the minimum eigenvalue of $X(\mathbf{x})$. We start with $L_{f}$ and find by using the bound provided by merikoski1997bounds
@@
The trace can be upper bounded by $\operatorname{Tr}(X)\leq mN\max_{i}\left\|\mathbf{v}_{i}\right\|_{2}^{2}$ and by knowing that $\log\det(X(\mathbf{x}))\leq-\log\det(X_{0})$, we find